OpenStudy (ittybitty14):

MEDAL MEDAL MEDAL MEDAL MEDAL MEDAL MEDAL!!!!!!!! Arrange these numbers from least to greatest: -5/8,-7/9,-4/5,-3/7

OpenStudy (ittybitty14):

explanation please! :)

OpenStudy (anonymous):

find the common denominator

OpenStudy (ittybitty14):

-7/9? I'm not sure!

OpenStudy (anonymous):

the least common denominator is the lowest number the denominators can all be multiplied into. in this case the least common denominator is 2520

OpenStudy (anonymous):

if you rewrite your numbers with the common denominator of 2520 it would look like this -1575/2520, -1960/2520, -2016/2520, -1080/2520

OpenStudy (anonymous):

this question is trying to trick you by making you label them from least to greatest

OpenStudy (ittybitty14):

ohhh okay! I think I'm starting to understand

OpenStudy (anonymous):

just to help out a little more: -1575/2520 = -5/8 -1960/2520 = -7/9 -2016/2520 = -4/5 -1080/2520 = -3/7

OpenStudy (anonymous):

can you order them least to greatest?

OpenStudy (ittybitty14):

I think -3/7, -4/5,-5/8, -7/9 but I'm not positive

OpenStudy (anonymous):

these numbers are negative so which number is furthest away from 0?

OpenStudy (ittybitty14):

-4/5?

OpenStudy (anonymous):

yes so that is the your lowest, you are labeling them from least to greatest. which comes next?

OpenStudy (ittybitty14):

-7/9,-5/8,-3/7

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

well done

OpenStudy (ittybitty14):

thanks!